# The Global Sections of Chiral de Rham Complexes on Closed Complex Curves

### v2

@@ 4. Global Sections on closed Complex Curves
By the inductive hypothesis ¯$\bar{\partial}^{\prime}a_{s-t+1}+F_{1}a_{s-t+2}+F_{2}a_{s-t+3}=0$, we have:
@@
¯¯$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{1}a_{s-t+1}$ $=-d\bar{\gamma}iH^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}(\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}+\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}}))F_{1}a_{s-t+1}$
@@
¯¯$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{1}a_{s-t+1}$ $=((t-1)(l-s)+\frac{1}{2}(t-1)(t-2))F_{1}a_{s-t+1}+(l-s+t-1)F_{1}a_{s-t+1}$
@@
¯¯$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{2}a_{s-t+2}$ $=-d\bar{\gamma}iH^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}(\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}+\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}}))F_{2}a_{s-t+2}$
@@
Moreover, by the inductive hypothesis ¯$\bar{\partial}^{\prime}a_{s-t+2}+F_{1}a_{s-t+3}+F_{2}a_{s-t+4}=0$, we have:
@@
¯¯$~{}~{}~{}\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{2}a_{s-t+2}$
@@
So under the inductive hypothesis, $a_{s-t}=\frac{-1}{t(l-s)+\frac{1}{2}t(t-1)}(\bar{\partial}^{\prime*}F_{1}a_{s-t+1}+\bar{\partial}^{\prime*}F_{2}a_{s-t+2})$ is a solution of the equation ¯$\bar{\partial}^{\prime}a_{s-t}+F_{1}a_{s-t+1}+F_{2}a_{s-t+2}=0$. ∎
So for any $a_{s}\in H^{0}(X,SW(\bar{T}^{*}X))[k,l,s]$ with $l-s>0$, we can give an $a=\sum_{i\geq 0}{a_{s-i}}$ satisfying the condition for global sections.
Case 3: $l-s=0$.
@@
$SW(\bar{T}^{*}X)[k,s,s]$ is a trivial vector bundle. So every holomorphic section $a_{s}$ of $SW(\bar{T}^{*}X)[k,s,s]$ satisfies $\nabla a=0$ and the space of holomorphic sections of $SW(\bar{T}^{*}X)[k,s,s]$ is isomorphic to its fibre $SW^{\gamma}[k,s,s]$.
@@
There exists a section $a_{s-1}$ of $SW(\bar{T}^{*}X)[k,s,s-1]$ such that ¯$\bar{\partial}^{\prime}a_{s-1}+F_{1}a_{s}=0$ if and only if $F_{1}a_{s}=0$.
@@
¯¯$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{1}a_{s}=(s-s)F_{1}a_{s}=0.$
@@
¯¯$\bar{\partial}^{\prime*}\bar{\partial}^{\prime}F_{1}a_{s}=0.$
@@
So ¯$\bar{\partial}^{\prime}a_{s-1}+F_{1}a_{s}=0$, if and only if
@@ 4.3. The $\mathfrak{sl}_{2}$ action on a $\beta\gamma-bc$ system
Let $V$ be a complex linear space of dimension one. Note that when the complex manifold $X$ is of dimension one, $\mathcal{W}^{\gamma}_{+}$ is isomorphic to $\mathcal{W}_{+}(V)$ as a vertex algebra by mapping $\bm{B}$, $\bm{\Gamma}$, $\bm{b}$, $\bm{c}$ to $\beta$, $\alpha$, $b$, $c$, respectively. Let $\tilde{\gamma}=\gamma-\gamma_{(-1)}=\sum_{n\neq-1}\gamma_{(n)}z^{-n-1}$, then the subalgebra of $\mathcal{W}_{+}(V)$ generated by $\bm{B}$, $\varGamma$, $\bm{b}$, $\bm{c}$ is isomorphic to the subalgebra of $\mathcal{W}_{+}$ generated by $\beta$, $\tilde{\gamma}$, $b$, $c$ by mapping $\bm{B}$, $\varGamma$, $\bm{b}$, $\bm{c}$ to $\beta$, $\tilde{\gamma}$, $b$, $c$, respectively.
Let $T$ be the Lie algebra generated by vector fields $\{x^{i}\frac{\partial}{\partial x}\}_{i=0,1,2}$ on $V$. $T$ is isomorphic to $\mathfrak{sl}_{2}$ with the following relations:
$[x^{i}\frac{\partial}{\partial x},x^{j}\frac{\partial}{\partial x}]=(j-i)x^{i+j-1}\frac{\partial}{\partial x}.$
@@
where $Q=:\beta c:$.
Therefore, $\mathcal{L}$ is an action of $\mathfrak{sl}_{2}$ on a $\beta\gamma-bc$ system $\mathcal{W}(V)$. And by the isomorphism between $\mathcal{W}^{\gamma}_{+}$ and $\mathcal{W}_{+}(V)$, $\mathcal{L}$ also induces a natural action of $\mathfrak{sl}_{2}$ on $SW^{\gamma}$.
Define $\mathcal{L}^{+}:T\rightarrow End_{\mathbb{C}}(\mathcal{W}(V))$ by:
$\mathcal{L}^{+}(x^{i}\frac{\partial}{\partial x})=(:\tilde{\gamma}^{i}\beta:)_{(0)}-i(::\tilde{\gamma}^{i-1}b:c:)_{(0)}.$
@@
The subspace of holomorphic sections of $\bigoplus_{k,s}SW(\bar{T}^{*}X)[k,s,s]$ with $F_{1}a_{s}=0$ is isomorphic to $\mathcal{W}^{T}(V)$.
@@
is equivalent to $F_{1}a_{s}=0$ by the isomorphism between $SW^{\gamma}$ and $\mathcal{W}_{+}(V)$. ∎
@@
We derive a description of the space of global sections of $\Omega^{ch}_{X}$ on the complex curve $X$ from the computations in the preceding part:
@@
If $X$ is a closed complex curve with genus $g\geq 2$, then the space of global sections:
(4.7) $H^{0}(X,\Omega^{ch}_{X})\cong\bigoplus_{k,s}((H^{0}(X,SW(\bar{T}^{*}X))[k,s,s]\cap\operatorname{Ker}F_{1})\bigoplus(\bigoplus_{l>s}H^{0}(X,SW(\bar{T}^{*}X))[k,l,s]))$
Moreover, the subspace $\bigoplus_{k,s}(H^{0}(X,SW(\bar{T}^{*}X))[k,s,s]\cap\operatorname{Ker}F_{1})$ is isomorphic to the space of invariants under the action of $\mathfrak{sl}_{2}$ on the $\beta\gamma-bc$ system $\mathcal{W}(V)$, where $V$ is a complex linear space of dimension one.

### v3

@@ 4. Global Sections on closed Complex Curves
By the inductive hypothesis $\bar{\partial}^{\prime}a_{s-t+1}+F_{1}a_{s-t+2}+F_{2}a_{s-t+3}=0$, we have:
@@
$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{1}a_{s-t+1}$ $=-d\bar{\gamma}iH^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}(\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}+\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}}))F_{1}a_{s-t+1}$
@@
$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{1}a_{s-t+1}$ $=((t-1)(l-s)+\frac{1}{2}(t-1)(t-2))F_{1}a_{s-t+1}+(l-s+t-1)F_{1}a_{s-t+1}$
@@
$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{2}a_{s-t+2}$ $=-d\bar{\gamma}iH^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}(\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}+\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}}))F_{2}a_{s-t+2}$
@@
Moreover, by the inductive hypothesis $\bar{\partial}^{\prime}a_{s-t+2}+F_{1}a_{s-t+3}+F_{2}a_{s-t+4}=0$, we have:
@@
$~~~\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{2}a_{s-t+2}$
@@
So under the inductive hypothesis, $a_{s-t}=\frac{-1}{t(l-s)+\frac{1}{2}t(t-1)}(\bar{\partial}^{\prime*}F_{1}a_{s-t+1}+\bar{\partial}^{\prime*}F_{2}a_{s-t+2})$ is a solution of the equation $\bar{\partial}^{\prime}a_{s-t}+F_{1}a_{s-t+1}+F_{2}a_{s-t+2}=0$. ∎
So for any $a_{s}\in H^{0}(X,SW(\bar{T}^{*}X))[k,l,s]$ with ls>0l-s>0, we can give an $a=\sum_{i\geq 0}{a_{s-i}}$ satisfying the condition for global sections.
Case 3: ls=0l-s=0.
@@
$SW(\bar{T}^{*}X)[k,s,s]$ is a trivial vector bundle. So every holomorphic section $a_{s}$ of $SW(\bar{T}^{*}X)[k,s,s]$ satisfies $\nabla a=0$ and the space of holomorphic sections of $SW(\bar{T}^{*}X)[k,s,s]$ is isomorphic to its fiber $SW^{\gamma}[k,s,s]$.
@@
There exists a section $a_{s-1}$ of $SW(\bar{T}^{*}X)[k,s,s-1]$ such that $\bar{\partial}^{\prime}a_{s-1}+F_{1}a_{s}=0$ if and only if $F_{1}a_{s}=0$.
@@
$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{1}a_{s}=(s-s)F_{1}a_{s}=0.$
@@
$\bar{\partial}^{\prime*}\bar{\partial}^{\prime}F_{1}a_{s}=0.$
@@
So $\bar{\partial}^{\prime}a_{s-1}+F_{1}a_{s}=0$, if and only if
@@ 4.3. The $\mathfrak{sl}_{2}$ action on a $\beta\gamma-bc$ system
Let $V$ be a complex linear space of dimension one. Note that when the complex manifold $X$ is of dimension one, $\mathcal{W}^{\gamma}_{+}$ is isomorphic to $\mathcal{W}_{+}(V)$ as a vertex algebra by mapping $\bm{B}$, $\bm{\Gamma}$, $\bm{b}$, $\bm{c}$ to $\beta$, $\alpha$, $b$, $c$, respectively. Let
$\tilde{\gamma}(z)=\gamma(z)-\gamma_{(-1)}=\sum_{n\neq-1}\gamma_{(n)}z^{-n-1},$
which is a formal distribution, but is not a field in a vertex algebra. We naturally correspond $\bm{B}$, $\varGamma$, $\bm{b}$, $\bm{c}$ to $\beta$, $\tilde{\gamma}$, $b$, $c$, respectively.
Let $T$ be the Lie algebra generated by vector fields $\{x^{i}\frac{\partial}{\partial x}\}_{i=0,1,2}$ on $V$. $T$ is isomorphic to $\mathfrak{sl}_{2}$ by the following relations:
$x^{i}\frac{\partial}{\partial x},x^{j}\frac{\partial}{\partial x}]=(j-i)x^{i+j-1}\frac{\partial}{\partial x}.$
@@
where $Q=:\beta c:$. $\mathcal{L}$ is naturally an action of $\mathfrak{sl}_{2}$ on the $\beta\gamma-bc$ system $\mathcal{W}(V)$, and the coefficients $(Q_{(0)}:\gamma^{i}b:)_{(n)}$ of the fields $Q_{(0)}:\gamma^{i}b:$ generate an affine Kac-Moody algebra $\hat{\mathfrak{sl}}_{2}$ at level 0 inside $\mathcal{W}(V)$.
Define a linear operator $\mathcal{L}^{+}:T\rightarrow End_{\mathbb{C}}(\mathcal{W}(V))$ as zero modes of formal distributions:
$\mathcal{L}^{+}(x^{i}\frac{\partial}{\partial x})=(:\tilde{\gamma}^{i}\beta:)_{(0)}-i(::\tilde{\gamma}^{i-1}b:c:)_{(0)},$
which is obtained by expanding $\mathcal{L}(x^{i}\frac{\partial}{\partial x})$ as infinite sums of products of modes of $\beta$, $\gamma$, $b$ and $c$, and throwing out all terms where the mode $\gamma_{(-1)}$ appears. Note that $\mathcal{L}^{+}$ is not a homomorphism of Lie algebras. The relation between $\mathcal{L}$ and $\mathcal{L}^{+}$ is as follows:
@@
Proposition 4.1.
The subspace $\mathcal{W}^{T}(V)$ is a sub vertex algebra of $\mathcal{W}(V)$.
Proof.
For any $A$, $B$ $\in$ $\mathcal{W}(V)$, we have
$\mathcal{L}(x^{i}\frac{\partial}{\partial x})(A_{(n)}B)=(\mathcal{L}(x^{i}\frac{\partial}{\partial x})A)_{(n)}B+A_{(n)}(\mathcal{L}(x^{i}\frac{\partial}{\partial x})B).$
Therefore, $\left\{A\in\mathcal{W}(V)|\mathcal{L}(x^{i}\frac{\partial}{\partial x})A=0\right\}$ are sub vertex algebras of $\mathcal{W}(V)$ and $\mathcal{W}^{T}(V)=\cap_{i=0,1,2}\left\{A\in\mathcal{W}(V)|\mathcal{L}(x^{i}\frac{\partial}{\partial x})A=0\right\}$ is a sub vertex algebra of $\mathcal{W}(V)$. ∎
Let
(4.7) $M_{1}:=\bigoplus_{k,s}(H^{0}(X,SW(\bar{T}^{*}X))[k,s,s]\cap\operatorname{Ker}F_{1}),$
where $F_{1}$ is a first order differential operator.
Proposition 4.2.
The subspace $M_{1}$ is a sub vertex algebra of $H^{0}(X,\Omega^{ch}_{X})$.
Proof.
By Lemma 4.4, one only need to check that the subspace $\bigoplus_{k,s}(SW^{\gamma}[k,s,s]\cap\operatorname{Ker}F_{1})$ of the fiber $SW^{\gamma}$ is a vertex algebra, which is obvious. ∎
@@
The subspace $M_{1}$ is isomorphic to $\mathcal{W}^{T}(V)$ as vertex algebras.
@@
is equivalent to $F_{1}a_{s}=0$ by the isomorphism between $\mathcal{W}_{+}^{\gamma}$ and $\mathcal{W}_{+}(V)$.
Since the isomorphism above between $M_{1}$ and $\mathcal{W}^{T}(V)$ is compatible with the vertex isomorphism between $\mathcal{W}_{+}^{\gamma}$ and $\mathcal{W}_{+}(V)$ by mapping $\bm{B}$, $\bm{\Gamma}$, $\bm{b}$, $\bm{c}$ to $\beta$, $\alpha$, $b$, $c$, respectively, we see $M_{1}$ and $\mathcal{W}^{T}(V)$ are in fact isomorphic as vertex algebras. ∎
Let $\mathcal{C}(V)$ be the subalgebra of $\mathcal{W}(V)$ generated by
$L(z)=:\beta(z)\partial\gamma(z):-:b(z)\partial{c}(z):,$
and
$G(z)=:b(z)\partial\gamma(z):.$
We have $Q_{(0)},G_{(1)}]=L_{(1)}$, with the definitions of $Q,L,G$ agree with Eq. (2.1). It is easy to check that $\mathcal{C}(V)\subset\mathcal{W}^{T}(V)$.
We conjecture that $\mathcal{C}(V)=\mathcal{W}^{T}(V)$.
The Virasoro field $\beta\partial\gamma:-:b\partial c:\in\mathcal{W}^{T}(V)$ corresponds naturally to the canonical Virasoro field in $H^{0}(X,\Omega_{X}^{ch})$, which exists for any complex manifold. The structure of $\mathcal{W}^{T}(V)$ is independent of the genus $g$ and the choice of complex structure by its construction.
@@
Let
(4.8) $M_{2}:=\bigoplus_{k}\bigoplus_{l>s}H^{0}(X,SW(\bar{T}^{*}X))[k,l,s],$
where $H^{0}(X,SW(\bar{T}^{*}X))[k,l,s]$ is the space of holomorphic sections of an antiholomorphic vector bundle $SW(\bar{T}^{*}X)[k,l,s]$.
Recall that $\mathcal{W}^{T}(V)$ is a sub vertex algebra of $\mathcal{W}(V)$ which consists of $\mathfrak{sl}_{2}$ invariants.
The subspace $M_{2}$ is naturally a module over $M_{1}$, where $M_{1}$ is defined by Eq. (4.7). And by the isomorphism of vertex algebras between $M_{1}$ and $\mathcal{W}^{T}(V)$, the subspace $M_{2}$ is a module over $\mathcal{W}^{T}(V)$. $M_{2}$ is not a vertex algebra.
We derive a description of the space of global sections of $\Omega^{ch}_{X}$ on the closed complex curve $X$ with genus $g\geq 2$ from the computations in the preceding parts:
@@
If $X$ is a closed complex curve with genus $g\geq 2$, define spaces $M_{1}$ and $M_{2}$ by Eq. (4.7) and (4.8), then the space of global sections:
(4.9) $H^{0}(X,\Omega^{ch}_{X})\cong M_{1}\oplus M_{2},$
where $M_{1}\cong\mathcal{W}^{T}(V)$ as vertex algebras. $M_{1}$ is a sub vertex algebra of $H^{0}(X,\Omega^{ch}_{X})$ and $M_{2}$ is a module over $\mathcal{W}^{T}(V)$.
Definition 4.2.
Let
$H^{0}(X,\Omega_{X}^{ch})[k,l]:=\left\{a\in H^{0}(X,\Omega^{ch}_{X})|L_{(1)}a=ka,J_{(0)}a=la\right\},$
$H^{0}(X,\Omega_{X}^{ch})[k,l,l]:=H^{0}(X,\Omega_{X}^{ch})[k,l]\cap H^{0}(X,SW(\bar{T}^{*}X))[k,l,l],$
$H^{0}(X,\Omega^{ch}_{X})[k,l,\leq s]:=\big\{a\in H^{0}(X,\Omega^{ch}_{X})[k,l]|\text{$a$ is a section of }\bigoplus_{i\geq 0}SW(\bar{T}^{*}X)[k,l,s-i]\big\},$
$H^{0}_{+}(X,\Omega^{ch}_{X})[k,l]:=H^{0}(X,\Omega^{ch}_{X})[k,l,\leq(l-1)].$
By Theorem 4.6:
$H^{0}(X,\Omega^{ch}_{X})[k,l]=H^{0}_{+}(X,\Omega^{ch}_{X})[k,l]\oplus H^{0}(X,\Omega^{ch}_{X})[k,l,l].$
We show a method of calculating $\dim H^{0}_{+}(X,\Omega^{ch}_{X})[k,l]$ based on Theorem 4.6.
For any $k\geq 0$ and $l$, $s$ such that ls>0l-s>0, we have
(4.10) $\dim H^{0}(X,\Omega^{ch}_{X})[k,l,\leq s]=\sum_{i\geq 0}\dim H^{0}(X,SW(\bar{T}^{*}X))[k,l,s-i],$
which follows from the exact sequences holding for any $s^{\prime}<l$:
$0\rightarrow H^{0}(X,\Omega^{ch}_{X})[k,l,\leq(s^{\prime}-1)]\rightarrow H^{0}(X,\Omega^{ch}_{X})[k,l,\leq s^{\prime}]\rightarrow H^{0}(X,SW(\bar{T}^{*}X))[k,l,s^{\prime}]\rightarrow 0.$
In particular,
(4.11) $\dim H^{0}_{+}(X,\Omega^{ch}_{X})[k,l]$ $=\dim H^{0}(X,\Omega^{ch}_{X})[k,l,\leq(l-1)]$
$=\sum_{i>0}\dim H^{0}(X,SW(\bar{T}^{*}X))[k,l,l-i],$
and $\dim H^{0}(X,SW(\bar{T}^{*}X))[k,l,l-i]$ is calculated by
$\dim H^{0}(X,SW(\bar{T}^{*}X))[k,l,s]=\dim H^{0}(X,\bar{T}X^{\otimes n}\otimes\bar{T}^{*}X^{\otimes m})\dim SW^{\gamma}[k,l,s],$
where nm=lsn-m=l-s.
In fact, using the Riemann-Roch formula (cf. [[7], IV, Theorem 1.3]), one can derive:
$\dim H^{0}(X,\bar{T}X^{\otimes n}\otimes\bar{T}^{*}X^{\otimes m})=\begin{cases}0,&n-m\leq-1,\\ 1,&n-m=0,\\ g,&n-m=1,\\ (n-m)(2g-2)+1-g,&n-m\geq 2.\end{cases}$
Let
(4.12) $H(q,y,z)$ $=\prod^{\infty}_{n=1}(1-q^{n-1}y)(1-q^{n}y^{-1})\frac{1}{(1+q^{n}z)}\frac{1}{(1+q^{n}z^{-1})}$
$=\sum_{m,m^{\prime}\in\mathbb{Z},n\geq 0}a_{n,m,m^{\prime}}(-y)^{m}(-z)^{m^{\prime}}q^{n},$
then $\dim SW^{\gamma}[k,l,s]=a_{k,l,s}$.
The dimensions of the first few weight spaces can be calculated directly.
If $k=0$, then
$\dim H^{0}(X,\Omega^{ch}_{X})[0,1]=g,$
and the dimensions of the other weight spaces are 0.
If $k=1$, then
$\dim H^{0}(X,\Omega^{ch}_{X})[1,0]$ $=g,$
$\dim H^{0}(X,\Omega^{ch}_{X})[1,1]$ $=4g-3,$
$\dim H^{0}(X,\Omega^{ch}_{X})[1,2]$ $=3g-3,$
and the dimensions of the other weight spaces are 0.
If $k=2$, then
$\dim H^{0}(X,\Omega^{ch}_{X})[2,-1]$ $=1,$
$\dim H^{0}(X,\Omega^{ch}_{X})[2,0]$ $=5g-2,$
$\dim H^{0}(X,\Omega^{ch}_{X})[2,1]$ $1,\displaystyle=4g-1$,
$\dim H^{0}(X,\Omega^{ch}_{X})[2,2]$ $=9g-8,$
and the dimensions of the other weight spaces are 0.
We have seen that the dimensions of weight spaces depend on the genus $g$.